OpenStudy (animalover02):

i will metal!!!!! What is the slope of the line that passes through the points (–1, 13) and (4, –2)? A.- 1/3 B.the slop is undienified C.-3 D.3

OpenStudy (animalover02):

need correct answer though.

OpenStudy (michele_laino):

hint: the requested slope is given by the subsequent computation: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 2 - 13}}{{4 - \left( { - 1} \right)}} = ...?\]

OpenStudy (animalover02):

i wont metal unless you tell me the answer

OpenStudy (michele_laino):

I'm sorry, I can't give direct answers

OpenStudy (animalover02):

that would be -15 over 5

OpenStudy (animalover02):

so a?